# FLIP: Towards Fine-grained Alignment between ID-based Models and Pretrained Language Models for CTR Prediction

### v3

@@ V-A1 Datasets
V-A2 Evaluation Metrics
We adopt two popular metrics to evaluate the performance of different methods on the CTR task, following [14, 17]. AUC (Area Under the ROC Curve) measures the probability that the CTR predictor will assign a higher score to a randomly chosen positive item than a randomly chosen negative item. A higher AUC indicates a better performance. Logloss is the binary cross-entropy loss. A smaller Logloss indicates a better performance. Notably, a slightly higher AUC or a lower Logloss (e.g., 0.001) can be considered as a significant improvement in CTR prediction [45, 5, 85].
V-A3 Baselines
The baseline methods can be mainly classified into three categories: (1) ID-based models, (2) PLM-based methods, and (3) methods that combine both ID-based models and PLMs.
(1) ID-based models that take tabular data as input:
  • AFM [18] extends FM by learning the weights of feature interactions via attentional networks.
  • PNN [13] is a product-based network that feeds the inner products of features embeddings into a DNN.
  • Wide&Deep [43] is a general learning framework that combines a wide network and a deep network to achieve the advantages of both.
  • DCN [44] improves Wide&Deep by replacing the wide part with a cross network to learn explicit high-order feature interactions.
  • DeepFM [5] is a extension of Wide&Deep by replacing the wide network with FM to explicitly learn second-order feature interactions.
  • xDeepFM [45] uses a compressed interaction network (CIN) to capture high-order feature interactions in a vector-wise way.
  • AutoInt [17] leverages self-attention networks to learn high-order features interactions.
  • DCNv2 [14] modifies DCN by proposing a more expressive cross network to better capture explicit feature interactions.
  • AFN [46] adopts a logarithmic transformation layer to learn adaptive order feature interactions.
(2) PLM-based methods that take textual data as input:
  • CTR-BERT [22] is a two-tower structure that feeds user information and item information into their respective BERT [19] to make CTR predictions.
  • P5 [23] transforms various recommendation tasks into text generation tasks by employing the pretrained language model T5 [58]. For a fair comparison, we use P5 for a single task only (i.e., CTR prediction) with multiple templates.
  • PTab [21] pretrains the BERT model on masked language modeling task and finetunes this BERT on downstream prediction tasks.
(3) Methods that combine ID-based models and PLMs:
  • CTRL [25] is currently a state-of-the-art method for combining ID-based models and PLMs through instance-level knowledge distillation.
TABLE II: The overall performance of different models on three datasets. The best result is given in bold, and the second-best value is underlined. We also use wavy underline to denote the best baseline performance. Rel.Impr denotes the relative AUC improvement rate of FLIP against each baseline. The symbol $\ast$” indicates statistically significant improvement of FLIP over the best baseline with $p$-value $<$ 0.001.
@@ V-A3 Baselines
AFM 0.8449 0.3950 2.04% 0.7946 0.5116 1.45% 0.7630 0.5160 2.06%
PNN 0.8546 0.3946 0.88% 0.7956 0.5131 1.32% 0.7725 0.5055 0.80%
Wide&Deep 0.8509 0.3957 1.32% 0.7951 0.5116 1.38% 0.7684 0.5090 1.34%
DCN 0.8509 0.4056 1.32% 0.7957 0.5108 1.31% 0.7693 0.5086 1.22%
DeepFM 0.8539 0.3905 0.96% 0.7947 0.5122 1.43% 0.7671 0.5138 1.51%
xDeepFM 0.8454 0.3934 1.98% 0.7953 0.5108 1.36% 0.7720 0.5079 0.87%
AFN 0.8525 0.3868 1.13% 0.7932 0.5139 1.63% 0.7654 0.5118 1.74%
AutoInt 0.8509 0.4013 1.32% 0.7953 0.5118 1.36% 0.7716 0.5071 0.92%
DCNv2 0.8548 0.3893 0.85% 0.7956 0.5103 1.32% 0.7724 0.5057 0.82%
CTR-BERT 0.8304 0.4131 3.82% 0.7795 0.5300 3.41% 0.7385 0.5316 5.44%
P5 0.8304 0.4173 3.82% 0.7801 0.5261 3.33% 0.7365 0.5336 5.73%
PTab 0.8426 0.4195 2.31% 0.7880 0.5384 2.30% 0.7456 0.5268 4.44%
CTRL 0.8572 0.3838 0.57% 0.7985 0.5101 0.95% 0.7741 0.5045 0.59%
@@
FLIP (Ours) 0.8621* 0.3788* - 0.8061* 0.5004* - 0.7787* 0.5001* -
V-A4 Implementation Details
All of our experiments are performed on 8 NVIDIA Tesla V100 GPUs with PyTorch [86]. In the modality alignment pretraining stage, we set the text and tabular mask ratio $r_{text}$ and $r_{tab}$ both to 15%. Unless specified otherwise, we adopt TinyBERT [87] as the basic PLM, and DCNv2 [14] as the basic ID-based model. During pretraining, the model is trained for 30 epochs with the AdamW [88] optimizer and a batch size of 1024. The learning rate is initialized as 5e-5 followed by a cosine decay strategy. The number of noise features $K$ is 25 in Eq. 10. The temperature $\tau$ is 0.7 in Eq. 13. In the supervised finetuning stage, we adopt the Adam [89] optimizer with learning rate selected from {1e-5,5e-5,1e-4,5e-4,1e-3,5e-3}. The finetuning batch size is 256 for MovieLens-1M and BookCrossing, and is 2048 for GoodReads.
V-B Performance Comparison (RQ1)
[…]
  • FLIP outperforms all baselines significantly and consistently. On MovieLens-1M, FLIP outperforms the best baseline by 0.005 AUC points. On BookCrossing, FLIP achieves up to 0.007 AUC improvements compared to the best baseline. On GoodReads, FLIP outperforms the best baseline by 0.004 AUC points. These results confirm the excellence of our proposed fine-grained alignment and jointly finetuning approach.
@@ V-B Performance Comparison (RQ1)
  • The baseline CTRL generally achieves the best performance among all the baselines. CTRL distills the PLM’s knowledge into an ID-based model and uses this ID-based model to make prediction, which ignores the PLM’s inherent capabilities, such as reasoning. Furthermore, it only focuses on instance-level alignment, resulting in learned representations with high similarity (Section V-G). Our FLIP combines ID-based models and PLMs through fine-grained alignment and model ensemble, effectively mitigating these limitations and improving performance.
V-C Compatibility Analysis (RQ2)
FLIP serves a model-agnostic framework that is compatible for various backbone models. Therefore, in this section, we investigate the model compatibility of FLIP in terms of both ID-based models and PLMs.
V-C1 Compatibility with ID-based models
@@
  • By tuning both the ID-based model and PLM, FLIP achieves the best performance on different backbone ID-based models significantly and consistently, demonstrating the superior compatibility of FLIP in terms of ID-based models.
TABLE III: The compatibility analysis of FLIP with different ID-based models on MovieLens-1M. The PLM is fixed as TinyBERT. N/A means to train the raw ID-based model from scratch. For each ID-based model, the best result is in bold, and the second-best is underlined. $\ast$” indicates statistically significant improvement over the raw ID-based model with $p$-value $<$ 0.001.
ID-based Model Finetuning Strategy MovieLens-1M
AUC Logloss
DeepFM N/A 0.8539 0.3905
@@
FLIP 0.8621* 0.3788*
TABLE IV: The compatibility analysis of FLIP with different ID-based models on BookCrossing. The PLM is fixed as TinyBERT. N/A means to train the raw ID-based model from scratch. For each ID-based model, the best result is in bold, and the second-best is underlined. $\ast$” indicates statistically significant improvement over the raw ID-based model with $p$-value $<$ 0.001.
ID-based Model Finetuning Strategy BookCrossing
AUC Logloss
DeepFM N/A 0.7947 0.5122
[…]
FLIP 0.8033* 0.5031*
AutoInt N/A 0.7943 0.5118
[…]
FLIP 0.8011* 0.5050*
DCNv2 N/A 0.7956 0.5103
[…]
FLIP 0.8061* 0.5004*
V-C2 Compatibility with PLMs
Similarly, we fix the ID-based model as DCNv2, and train FLIP with three different PLMs, including TinyBERT with 14.5M parameters, RoBERTa-Base [20] with 125M parameters, and RoBERTa-Large [20] with 355M parameters. The results are summarized in Table V and  VI, from which we observe that:
[…]
  • Although a larger language model is expected to bring about more significant performance improvements, we find that such performance gains might gradually diminishes. Due to the page limitation and resource constraint, we leave the further analysis on such a phenomenon as future works (e.g., involving further larger PLMs).
TABLE V: The compatibility analysis of FLIP with different PLMs on MovieLens-1M. The ID-based model is fixed as DCNv2. For each type of PLM, the best result is in bold, and the second-best is underlined. $\ast$” indicates statistically significant improvement over raw DCNv2 model with $p$-value $<$ 0.001.
PLM Finetuning Strategy MovieLens-1M
AUC Logloss
DCNv2 0.8548 0.3893
[…]
FLIP 0.8621* 0.3788*
[…]
FLIP 0.8634* 0.3770*
[…]
FLIP 0.8650* 0.3757*
TABLE VI: The compatibility analysis of FLIP with different PLMs on BookCrossing. The ID-based model is fixed as DCNv2. For each type of PLM, the best result is in bold, and the second-best is underlined. $\ast$” indicates statistically significant improvement over raw DCNv2 model with $p$-value $<$ 0.001.
PLM Finetuning Strategy BookCrossing
AUC Logloss
DCNv2 0.7956 0.5103
[…]
FLIP 0.8061* 0.5004*
[…]
FLIP 0.8083* 0.4995*
[…]
FLIP 0.8092* 0.4986*
Figure 6: The hyperparameter study on textual mask ratio $r_{text}$ and tabular mask ratio $r_{tab}$ on MovieLens-1M (left two images) and BookCrossing (right two images).
V-D Ablation Study (RQ3)
We conduct comprehensive ablation experiments for better understanding the contributions of different components in our proposed FLIP. In this section, all the experiments are conducted on Movielens-1M and BookCrossing datasets.
Firstly, we evaluate the impact of pretraining objectives (i.e., MLM, MTM, ICL) by eliminating them from the pretraining stage. Note that removing all three objectives means that the pretraining stage does not exist. The results are reported in Table VII.
  • As we can observe, the optimal performance is achieved when three losses are deployed simultaneously, and removing each loss will degrade performance, while eliminating all losses results in the lowest performance. These phenomena demonstrate every component contributes to the final performance.
  • Removing ICL (w/o ICL) obtains better performance than removing MLM&MTM (w/o MLM&MTM), indicating that joint modeling for MLM&MTM tasks can learn and align meaningful representations even without ICL.
Next, to further investigate the effect of joint reconstruction for MLM and MTM, we design the following two variants:
  • w/o Field-level Masking: We replace the field-level masking for textual data with the common random token-level masking, as discussed in Section IV-C1.
  • w/o Joint Reconstruction: The reconstruction of one masked modality no longer relies on the help from the other modality.
The results are shown in Table VII. The performance drops, especially on the Logloss metric, when we either remove the field-level masking strategy or eliminate the joint reconstruction. Such a phenomenon demonstrates the importance of fine-grained feature-level alignment, which ensures the feature-level communication and interaction between ID-based models and PLMs for dual modalities.
TABLE VII: The results of ablation study. The best values are given in bold.
@@ V-D Ablation Study (RQ3)
Figure 7: The hyperparameter study on the temperature $\tau$.
(a) FLIPw/oMLM&MTM
(b) FLIPw/oMLM
(c) FLIPw/oMTM
@@
Figure 8: The visualization of feature representations learned by different model variants on MovieLens-1M dataset. We use SVD to project the feature embedding matrix into 2D data. Note that FLIPw/oMLM&MTM indicates the instance-level contrastive learning (ICL) variant without the MLM and MTM pretraining objectives.
Figure 9: The visualization of similarities between the sample representations of masked textual and tabular data. “Text-$f$” and “Tab-$f$” denote that we mask the $f$-th field of the input data of textual or tabular modalities, respectively.
V-E Hyperparameter Study (RQ4)
In this section, we conduct experiments to study the influence of different hyperparameters on the FLIP.
V-E1 The Impact of Mask Ratio $r_{text}$ and $r_{tab}$
Since there are two independent mask ratio for textual and tabular modalities, we fix the mask ratio on the one side and alter the mask ratio on the other side from $\{0.15,0.3,0.6\}$. The results are in Figure 6, from which we observe that the best performance is generally achieved when both mask ratios are relatively small (i.e., 0.15). The reason is that excessive masking might lead to ambiguity in the target modality data, which hurts the model pretraining [19, 20].
V-E2 The Impact of Temperature Hyperparameter $\tau$
The temperature $\tau$ controls the sharpness of estimated distribution. We select $\tau$ from $\{0.05,0.1,0.3,0.5,0.7,1.0\}$, and report the results in Figure 7. When the temperature gradually grows, the performance increases first and then decreases, with the optimal temperature choice between $0.5$ and $0.7$. As suggested in previous works [90, 91], a too small temperature would only concentrate on the nearest sample pairs with top scores, and a too large temperature might lead to a flat distribution where all negative sample pairs receive almost the same amount of punishment. Both of them will hurt the final performance [90, 25]. Therefore, we choose $\tau=0.7$ in our approach.
V-F Analysis on Feature-level Alignment (RQ5)
We conduct case studies on Movielens-1M and BookCrossing datasets to further explore the fine-grained feature-level alignment established during pretraining. For a dual-modality input pair $(\textbf{x}_{i}^{text},\textbf{x}_{i}^{tab})$, we first perform field-level data masking to mask out each field respectively, resulting in $F$ pairs of corrupted inputs $\{(\textbf{x}_{i,(f)}^{text},\textbf{x}_{i,(f)}^{tab})\}_{f=1}^{F}$, where $\textbf{x}_{i,(f)}^{text}$ and $\textbf{x}_{i,(f)}^{tab}$ denote that we solely mask the $f$-field of the textual or tabular data. Then we employ the PLM and ID-based model to encode them into $F$ pairs of normalized sample representations $\{(\textbf{z}_{i,(f)}^{text},\textbf{z}_{i,(f)}^{tab})\}_{f=1}^{F}$. Next, we compute the mutual similarity scores (measured by dot product) over each cross-modal representation pair, and visualize the heat map in Figure 9.
V-G Quality of Feature Representations (RQ6)

### v4

@@ 4.1.3. Baselines
Table 2. The overall performance of different models from three categories (i.e., ID-based, PLM-based, and ID+PLM). For each type of models, the best result is given in bold, and the second-best value is underlined. Rel.Impr denotes the relative AUC improvement rate of our method against each baseline within each category. The symbol $\ast$” indicates statistically significant improvement of FLIP over the best baseline with $p$-value $<$ 0.001.
@@
ID-based AFM 0.8449 0.3950 1.79% 0.7946 0.5116 1.06% 0.7630 0.5160 1.96%
PNN 0.8546 0.3946 0.63% 0.7956 0.5131 0.93% 0.7725 0.5055 0.70%
Wide&Deep 0.8509 0.3957 1.07% 0.7951 0.5116 1.00% 0.7684 0.5090 1.24%
DCN 0.8509 0.4056 1.07% 0.7957 0.5108 0.92% 0.7693 0.5086 1.12%
DeepFM 0.8539 0.3905 0.71% 0.7947 0.5122 1.04% 0.7671 0.5138 1.41%
xDeepFM 0.8454 0.3934 1.72% 0.7953 0.5108 0.97% 0.7720 0.5079 0.77%
AFN 0.8525 0.3868 0.88% 0.7932 0.5139 1.24% 0.7654 0.5118 1.64%
AutoInt 0.8509 0.4013 1.07% 0.7953 0.5118 0.97% 0.7716 0.5071 0.82%
DCNv2 0.8548 0.3893 0.61% 0.7956 0.5103 0.93% 0.7724 0.5057 0.72%
[…]
PLM-based CTR-BERT 0.8304 0.4131 1.88% 0.7795 0.5300 1.65% 0.7385 0.5316 1.23%
P5 0.8304 0.4173 1.88% 0.7801 0.5261 1.58% 0.7365 0.5336 1.51%
PTab 0.8426 0.4195 0.41% 0.7880 0.5384 0.56% 0.7456 0.5268 0.27%
[…]
ID+PLM CTRL 0.8572 0.3838 0.57% 0.7985 0.5101 0.95% 0.7741 0.5045 0.59%
MoRec 0.8561 0.3896 0.70% 0.7990 0.5087 0.89% 0.7731 0.5085 0.72%
FLIP (Ours) 0.8621* 0.3788* - 0.8061* 0.5004* - 0.7787* 0.5001* -
4.1.4. Implementation Details
All of our experiments are performed on 8 NVIDIA Tesla V100 GPUs with PyTorch (Paszke et al., 2019). In the modality alignment pretraining stage, we set the text and tabular mask ratio $r_{text}$ and $r_{tab}$ both to 15%. Unless specified otherwise, we adopt TinyBERT (Jiao et al., 2020) as the PLM, and DCNv2 (Wang et al., 2021) as the ID-based model. During pretraining, the model is trained for 30 epochs with the AdamW (Loshchilov and Hutter, 2017) optimizer and a batch size of 1024. The learning rate is initialized as 5e-5 followed by a cosine decay strategy. The number of noise features $K$ is 25 in Eq. 10. The temperature $\tau$ is 0.7 in Eq. 11. In the adaptive finetuning stage, we adopt the Adam (Kingma and Ba, 2014) optimizer with learning rate selected from {1e-5,5e-5,1e-4,5e-4,1e-3,5e-3}. The finetuning batch size is 256 for MovieLens-1M and BookCrossing, and is 2048 for GoodReads.
4.2. Performance Comparison
[…]
  • FLIP outperforms all baselines from all three categories (i.e., ID-based, PLM-based, and ID+PLM) significantly, which confirms the excellence of our proposed fine-grained feature-level alignment and adaptive finetuning approach for CTR prediction.
[…]
  • CTRL and MoRec generally outperform other baselines due to their integration of ID-based models and PLMs. However, they either overlook the fine-grained alignment between IDs and tokens or ignore cross-modal interactions, thereby degrading the performance. FLIP addresses these shortcomings by bridging ID-based models and PLMs through jointly masked tabular/language modeling, which enables fine-grained feature-level interactions between dual modalities, thus resulting in superior performance.
4.3. Compatibility Analysis
FLIP serves as a model-agnostic framework that is compatible with various backbone models. In this section, we investigate the model compatibility in terms of different ID-based models and PLMs.
4.3.1. Compatibility with ID-based models
We apply FLIP to three different ID-based models, including DeepFM, AutoInt and DCNv2, while keeping TinyBERT as the PLM. The results are listed in Table 3. We can obtain the following observations:
[…]
  • By jointly tuning the ID-based model and PLM, FLIP achieves the best performance across various ID-based backbone models significantly, demonstrating the superior compatibility of FLIP in terms of ID-based models.
Table 3. The compatibility w.r.t. different ID-based models. The PLM is fixed as TinyBERT. N/A means to train the vanilla ID-based model from scratch. For each ID-based model, the best result is in bold, and the second-best is underlined.
ID-based Model Finetuning Strategy MovieLens-1M BookCrossing
AUC Logloss AUC Logloss
DeepFM N/A 0.8539 0.3905 0.7947 0.5122
@@ 4.3.1. Compatibility with ID-based models
FLIP 0.8615 0.3758 0.8033 0.5031
AutoInt N/A 0.8509 0.4013 0.7943 0.5118
@@
FLIP 0.8600 0.3807 0.8011 0.5050
DCNv2 N/A 0.8548 0.3893 0.7956 0.5103
@@
FLIP 0.8621 0.3788 0.8061 0.5004
4.3.2. Compatibility with PLMs
Similarly, we keep DCNv2 as the ID-based model, and select PLMs of different sizes, including TinyBERT (14.5M), RoBERTa-Base (125M) (Liu et al., 2019a) , and RoBERTa-Large (355M) (Liu et al., 2019a). The results are in Table 4, from which we find that:
[…]
  • While increasing the PLM’s size is expected to yield more notable performance improvements, the advantages of scaling up gradually taper off. For instance, the improvement from RoBERTa-Base to RoBERTa-large is significantly smaller than the improvement from TinyBERT to RoBERTa-Base.
[…]
Table 4. The compatibility w.r.t. different PLMs. The ID-based model is fixed as DCNv2. For each type of PLM, the best result is in bold, and the second-best is underlined.
PLM Finetuning Strategy MovieLens-1M BookCrossing
AUC Logloss AUC Logloss
DCNv2 0.8548 0.3893 0.7956 0.5103
@@ 4.3.2. Compatibility with PLMs
FLIP 0.8650 0.3757 0.8092 0.4986
4.4. Ablation Study
We conduct ablation experiments for better understanding the contributions of different components in our proposed FLIP.
Table 5. The results of ablation study.
@@ 4.4. Ablation Study
Firstly, we evaluate the impact of pretraining objectives (i.e., MLM, MTM, ICL) by eliminating them from the pretraining stage. Note that removing all three objectives means that the pretraining stage does not exist. The results are reported in Table 5.
  • As we can observe, the optimal performance is achieved when three losses are deployed simultaneously, and removing each loss will degrade performance, while eliminating all losses results in the lowest performance. These phenomena demonstrate that each component contributes to the final performance.
  • Removing ICL (w/o ICL) obtains better performance than removing MLM&MTM (w/o MLM&MTM), indicating that joint modeling for MLM&MTM tasks can learn meaningful cross-modal alignments even without ICL.
Next, to further investigate the effect of jointly masked modality modeling, we design the following two variants:
  • w/o Field-level Masking: We replace the field-level masking for textual data with the common random token-level masking, as discussed in Section 3.3.1.
  • w/o Joint Reconstruction: The reconstruction of one masked modality depends only on itself and no longer relies on the help from the other modality.
The results are shown in Table 5. The performance drops, especially on the Logloss metric, when we either remove the field-level masking strategy or eliminate the joint reconstruction. Such a phenomenon demonstrates the importance of fine-grained feature-level alignment, which ensures the feature-level communication and interaction between ID-based models and PLMs for dual modalities.
4.5. Hyperparameter Study
Figure 3. The hyperparameter study on textual mask ratio $r_{text}$ (left column) and tabular mask ratio $r_{tab}$ (right column) on MovieLens-1M (top) and BookCrossing (bottom) datasets.
4.5.1. The Impact of Mask Ratio $r_{text}$ and $r_{tab}$
Since there are two independent mask ratio for textual and tabular modalities, we fix the mask ratio on the one side and alter the mask ratio on the other side from $\{0.15,0.3,0.6\}$. The results are in Figure 3, from which we observe that the best performance is generally achieved when both mask ratios are relatively small (i.e., 0.15). The reason is that excessive masking might lead to ambiguity in the target modality data, which hurts the model pretraining (Devlin et al., 2018; Liu et al., 2019a). So we set the $r_{text}$ and $r_{tab}$ both to 15%.
Figure 4. The hyperparameter study on the temperature $\tau$.
4.5.2. The Impact of Temperature Hyperparameter $\tau$
The temperature $\tau$ controls the sharpness of estimated distribution. We select $\tau$ from $\{0.05,0.1,0.3,0.5,0.7,1.0\}$, and report the results in Figure 4. When the temperature gradually grows, the performance increases first and then decreases, with the optimal temperature choice between $0.5$ and $0.7$. As suggested in previous works (Wang and Liu, 2021; Chen et al., 2020), a too small temperature would only concentrate on the nearest sample pairs with top scores, and a too large temperature might lead to a flat distribution where all negative sample pairs receive almost the same amount of punishment. Both of them will hurt the final performance (Wang and Liu, 2021; Li et al., 2023a). Therefore, we choose $\tau=0.7$ in our approach.
4.6. Analysis on Fine-grained Alignment
4.6.1. Feature-level Alignment
Figure 5. Visualization of similarities between the sample representations of masked textual and tabular data. ”Text-$f$” and ”Tab-$f$” denote that we mask the $f$-th field of the input data of textual or tabular modalities, respectively.
We conduct case studies to further explore the fine-grained feature-level alignment established during pretraining. For a dual-modality input pair $(\textbf{x}_{i}^{text},\textbf{x}_{i}^{tab})$, we first perform field-level data masking to mask out each field respectively, resulting in $F$ pairs of corrupted inputs $\{(\textbf{x}_{i,(f)}^{text},\textbf{x}_{i,(f)}^{tab})\}_{f=1}^{F}$, where $\textbf{x}_{i,(f)}^{text}$ and $\textbf{x}_{i,(f)}^{tab}$ denote that we solely mask the $f$-field of the textual or tabular data. Then we employ the PLM and ID-based model to encode them into $F$ pairs of normalized sample representations $\{(\textbf{z}_{i,(f)}^{text},\textbf{z}_{i,(f)}^{tab})\}_{f=1}^{F}$. Next, we compute the mutual similarity scores (measured by dot product) over each cross-modal representation pair, and visualize the heat map in Figure 5.
(a) FLIPw/o ${}_{\text{w/o }MLM\&MTM}$
(b) FLIPw/o ${}_{\text{w/o }MLM}$
(c) FLIPw/o ${}_{\text{w/o }MTM}$
@@ 4.6.1. Feature-level Alignment
Figure 6. The visualization of feature ID embeddings learned by different model variants on MovieLens-1M. We use SVD to project the feature embedding matrix into 2D data.
4.6.2. Visualization